# Causal Data Fusion for Panel Data without Pre-Intervention Period

### v1

@@ 6 Conclusion
We introduced two data fusion-based methods for identifying and estimating causal effects in panel data settings where no pre-intervention data are available. We introduced the equi-confounding data fusion and synthetic control data fusion methods, through which we addressed the challenges posed by sudden events that leave researchers without the typical pre-intervention period required by conventional causal inference panel data methods. We showed the efficacy of these methods by obtaining non-trivial bounds on the absolute biases, and provided conditions under which the bounds converge to zero. We showed that the synthetic control data fusion method is more flexible as it requires generally weaker assumptions. On the other hand, the equi-confounding methods are simpler to implement and to interpret. Our simulation results confirm our theoretical findings and demonstrate the fast decreasing rate of biases. Our causal panel data fusion methods have high potentials to be used in situations where sudden shocks preclude the collection of pre-intervention data. Future research could explore the applications of these methods in real-world cases, such as public health crises or epidemiological shocks.
@@ Acknowledgements
We would like to express our gratitude to the Massachusetts Department of Public Health (DPH). We extend special thanks to Boudu Bingay, Elizabeth T. Russo, and Joshua Norville, for their support throughout this process.

### v2

@@ 5.1 Placebo Test
Synthetic Covariates Chelsea Chelsea Average of 19 controls Median income - Hispanic 0.522 0.388 0.367 Median income - Black 0.206 0.371 0.388 Proportion - Hispanic 0.784 0.479 0.294 Proportion - Black 0.362 0.350 0.310 Median age - Hispanic 0.73 0.526 0.396 Median age - Black 0.418 0.414 0.406 Proportion aged 65+ - Hispanic 0.637 0.348 0.265 Proportion aged 65+ - Black 0.230 0.309 0.235
Table 1: Covariate means for Chelsea, synthetic Chelsea, and average of control units.
6 Application to COVID-19 Vaccination Rate Analysis
During the COVID-19 pandemic, community organizations played a pivotal role in promoting health equity through localized efforts. In our empirical study, we evaluated the effect of the pro-vaccination initiatives of La Colaborativa, a major Hispanic community-based organization in Chelsea, Massachusetts, on the COVID-19 vaccination rate of the Hispanic subpopulation of that city. Chelsea was among the 20 municipalities in Massachusetts most severely impacted by COVID-19. This disproportionate impact was evident in both case rates and social determinants of health, particularly among Hispanic (which comprises $60\%$ of Chelsea’s population) and Black sub-populations.
Despite the urgent need for vaccination, access barriers in the early rollout disproportionately affected non-English-speaking and low-income populations. When access first became available in Massachusets, eligibility alone did not guarantee access–many residents faced challenges such as language barriers, lack of reliable transportation, and difficulties securing online appointments. These disparities created an urgent need for targeted community-based interventions to ensure equitable vaccine distribution. To address this challenge, La Colaborativa launched a series of pro-vaccination efforts in March 2021, in alignment with the Massachusetts Vaccine Equity Initiative. These initiatives included door-to-door outreach, mobile vaccination clinics, and strategic partnerships with the East Boston Neighborhood Health Center to increase accessibility for Chelsea’s Hispanic population.
Evaluating the impact of these interventions poses a significant challenge due to the timing of the available data. The Massachusetts Public Health Department began recording COVID-19 vaccination rates in March 2021-coinciding with the start of La Colaborativa’s efforts. Hence, no pre-intervention data were collected. This lack of a pre-intervention period prevents the use of traditional causal inference approaches, which typically rely on pre-treatment trends to establish counterfactual comparisons. To address this limitation, we applied our proposed methodology to estimate the causal effects of La Colaborative’s intervention.
We considered the Hispanic sub-population in 20 cities in Massachusetts, as our target domain with monthly COVID-19 vaccination rate from March 2021 to July 2022 as the outcome variable, and with covariates median income, proportion of the subpopulation, median age, and proportion of residents aged 65 or older of the Hispanic sub-population. The city of Chelsea, which is our target unit, is among these 20 cities. The rest of the 20 cities were chosen based on similarities in their covariates to Chelsea. Importantly, cities served by similar community organizations to La Colaborativa were excluded from the donor pool to ensure a valid comparison group. We considered the Black sub-population in the same 20 cities as our reference domain. The reasoning behind this choice for the reference domain was that both Black and Hispanic populations face systemic barriers to healthcare access, vaccine hesitancy, and similar structural inequalities. These shared factors make Black vaccination trends a valuable reference for estimating causal effects in the Hispanic population.
Unit Weight Unit Weight
Bellingham 0.000 Marlborough 0.000
Chicopee 0.000 Methuen 0.000
Everett 0.228 Milford 0.000
Fitchburg 0.000 New Bedford 0.000
Haverhill 0.000 Peabody 0.000
Holyoke 0.310 Revere 0.330
Lawrence 0.037 Southbridge 0.000
Leominster 0.000 Springfield 0.094
Lynn 0.000 Waltham 0.000
Worcester 0.000
Table 2: Unit weights in the synthetic Chelsea.
(a) COVID-19 vaccination rates for Black sub-population (reference domain).
(b) COVID-19 vaccination rates for Hispanic sub-population (target domain).
Figure 6: COVID-19 vaccination rates for the observed units and synthetic Chelsea in reference and target domains.
(a) Placebo test for COVID-19 vaccination rates for Black sub-population (reference domain).
(b) Placebo test for COVID-19 vaccination rates for Hispanic sub-population (target domain).
Figure 7: Placebo test for COVID-19 vaccination rate analysis.
Results:
We used the value 0.1 for both $\eta_{Z}$ and $\eta_{X}$. However, as in the simulation study, we performed sensitivity analysis to investigate the sensitivity of our estimated causal effect to the choice of $\eta_{Z}$ and $\eta_{X}$, and we observed small impact of deviating these hyper-parameters (see Figure 5(b)). Table 1 presents the covariates used in the estimation process. As it can be seen, the method successfully generated the synthetic Chelsea with covariates close to those of the real Chelsea. Table 2 displays the weights assigned to each control city in constructing synthetic Chelsea. The results indicate that the best matching for Chelsea is achieved by combining Revere, Holyoke, Everett, Lawrence, and Springfield. All other units in the donor pool were assigned zero weights.
Figure 6 presents the COVID-19 vaccination rates for the observed units and synthetic Chelsea in reference and target domains. In Figure 6(a), synthetic Chelsea closely follows the actual Chelsea in the reference domain, demonstrating that our algorithms has succeeded to obtain a good match. This, along with the high degree of balance in observed covariates, supports the validity of synthetic Chelsea as a counterfactual for actual Chelsea in the absence of the intervention. In Figure 6(b), the observed Chelsea Hispanic COVID-19 vaccination rate and its synthetic version begin to diverge after the intervention starts, with the gap widening over time. Specifically, it can be seen that Chelsea has a steeper rise in the earlier months. The divergence between the black solid line (actual Chelsea) and the black dashed line (synthetic Chelsea) suggests that La Colaborativa’s efforts contributed to a greater increase in Hispanic vaccination rates than would have been expected without the intervention. We have also provided the results for placebo test in Figure 7(a), which confirms the statistical significance of our findings regarding the causal effect.
We now present the estimated causal effect of the community organization intervention based on our proposed methodologies. Over the 16-month period from March 2021 to July 2022, the COVID-19 vaccination rates for Hispanic sub-population was estimated to be increased by $13.6\%$ using the linear equi-confounding method, $13.2\%$ using the logarithmic equi-confounding method, and $10.1\%$ using the synthetic control data fusion method. All three methods indicate a similar large positive causal effect, which affirms the effectiveness of the community organization intervention.
@@ 7 Conclusion
We introduced two data fusion-based methods for identifying and estimating causal effects in panel data settings where no pre-intervention data are available. We introduced the equi-confounding data fusion and synthetic control data fusion methods, through which we addressed the challenges posed by sudden events that leave researchers without the typical pre-intervention period required by conventional causal inference panel data methods. We showed the efficacy of these methods by obtaining bounds on the biases, and provided conditions under which the bounds converge to zero. We showed that the synthetic control data fusion method is more flexible as it requires generally weaker assumptions. On the other hand, the equi-confounding methods are simpler to implement and to interpret. Our simulation results confirm our theoretical findings and demonstrate the fast decreasing rate of the biases. Our proposed causal panel data fusion methods have high potentials to be used in situations where sudden shocks preclude the collection of pre-intervention data. As an example of such applications, we applied our proposed methods to estimate the causal effect of a community organization vaccination intervention in Chelsea, Massachusetts. Our results demonstrate that over the 16-month period from March 2021 to July 2022, the Hispanic COVID-19 vaccination rate in Chelsea increased by $13.6\%$ using the linear equi-confounding method, $13.2\%$ using the logarithmic equi-confounding method, and $10.1\%$ using the synthetic control data fusion method. All three methods indicate a similar large positive causal effect, which affirms the effectiveness of the community organization intervention.
@@ Acknowledgements
We would like to express our gratitude to the Massachusetts Department of Public Health (DPH). We extend special thanks to Boudu Bingay, Elizabeth T. Russo, and Joshua Norville, for their support throughout this project.